mikewwe13:

Solve for f. d = 16ef² A. f = ± √de/4e B. f = ± 4√de/e C. f = ± 4√de D. f = ± √de/16

mikewwe13:

@Vocaloid

Vocaloid:

try to think about what steps you need to take to isolate f for example, since f is being multiplied by 16 and e what do you need to do to "undo" this?

mikewwe13:

well the answer is A correct ?

Vocaloid:

yeah that's it

mikewwe13:

Solve for e. T = e^2/9 A. e = ± 3√T ​​B. e = ± 3√T/T C.​ e = ± 9√T​ D. e = ±3T

Vocaloid:

any ideas?

mikewwe13:

the answer is C or B

mikewwe13:

the answer is B

Vocaloid:

T = e^2/9 start by multiplying both sides by 9, what do you get?

mikewwe13:

81

Vocaloid:

e^2/9 * 9 = ?

mikewwe13:

73

Vocaloid:

hint: notice how e^2 is being divided by 9 and then multiplied by 9 so they cancel out

mikewwe13:

ok that's true

Vocaloid:

so e^2 / 9 * 9 = e^2 e^2 = 9T solve for e

mikewwe13:

= 0

mikewwe13:

i'm still working on my calculations

Vocaloid:

hint: you have e^2 on one side so simply take the square root of both sides to isolate e e = sqrt(9T) and then simplify until it looks like one of the answer choices

mikewwe13:

T = e/9

Vocaloid:

what's the square root of 9?

mikewwe13:

3

mikewwe13:

OH I see

Vocaloid:

good, so sqrt(9T) = +/- 3sqrt(T) answer choice A

mikewwe13:

it's A

Vocaloid:

yes

mikewwe13:

The formula represents the distance x traveled by an object with acceleration a for t seconds. x = 1/2at^2 Solve for t. A. t = ±√x/2a B. t = ± 2ax^2 C. t = ± √2xa/a D. t = ± 1/2√ax

mikewwe13:

THE ANSWER IS C

Vocaloid:

yeah good

mikewwe13:

The formula gives the volume V of a right cylinder with radius r and height h. V=πr^2h Solve for r. A. r = √Vπh B. r = π√Vh C. r = √Vπh/h D. r = √Vπh/πh

mikewwe13:

THE ANSWER IS D

Vocaloid:

good

mikewwe13:

Solve for t. d = −16t^2 + 8t A. t = 1/4 ± 8√1 − d​ B. t = 8 ± √1 − d C. t = 1/2 ± 8√1 − d D. t = 1/4 ± √1−d/4

mikewwe13:

The Answer is D

mikewwe13:

is it correct cause i'm not sure

Vocaloid:

good

mikewwe13:

D is correct ?

Vocaloid:

yes

mikewwe13:

thank you Vocaloid once again for your help